0 = 6 + 3x − 2y find the slope
Answer by Timnewman(323)

0 = 6 + 3x − 2y find the slope
sol.
3x-2y+6=0
put the equation in the form of:
y=mx+c where m is the slope.
Therefore,
-2y=-3x-6
divide through by -2
y=3x/2+3.
The slope is 3/2